# 3D-GRAND: Towards Better Grounding and Less Hallucination for 3D-LLMs

### v1

@@ 6 Conclusion
In this paper, we introduced 3D-GRAND, a large-scale, densely-grounded 3D-text dataset designed for grounded 3D instruction tuning, and 3D-POPE, a comprehensive benchmark for evaluating object hallucination in 3D-LLMs. Through extensive experiments, we demonstrated the effectiveness of our 3D-LLM in improving grounding and reducing hallucination, achieving state-of-the-art performance on the ScanRefer and 3D-POPE benchmarks. Our ablation study and qualitative analysis highlighted the importance of large-scale densely-grounded instruction tuning and the choice of pre-trained language models in developing high-performing 3D-LLMs. We believe that our contributions will inspire further research and innovation in this field, ultimately leading to the development of more advanced and capable 3D-LLMs for a wide range of applications.
@@ A.1 Model input and output demonstration
Appendix B 3D-GRAND Data Curation
@@ D.1.2 Crowd-sourcing Validity
Before presenting any tasks to the workers, we present them with a set of instruction tasks that show an example annotation, the correct response (as determined by us), and the reason why that response is correct. They are paired with an incorrect example and an explanation of why it is incorrect in order to ensure unbiased annotations. Examples of qualifying instructions are shown in 10. These instructions are intentionally brief, as we found through trial-and-error that longer, paragraph-based instructions were largely ignored by annotators.

### v2

# 3D-GRAND: A Million-Scale Dataset for 3D-LLMs with Better Grounding and Less Hallucination

@@ 6 Conclusion
In this paper, we introduced 3D-GRAND, a large-scale, densely-grounded 3D-text dataset designed for grounded 3D instruction tuning, and 3D-POPE, a comprehensive benchmark for evaluating object hallucination in 3D-LLMs. Through extensive experiments, we demonstrated the effectiveness of our dataset on 3D-LLMs in improving grounding and reducing hallucination, achieving state-of-the-art performance on the ScanRefer and 3D-POPE benchmarks. Our ablation study and qualitative analysis highlighted the importance of densely-grounded instruction tuning, the data scaling law, and effective sim-to-real transfer in developing high-performing 3D-LLMs. We hope our contributions and findings can spark further research and innovation in this field, ultimately leading to the development of more advanced and capable 3D-LLMs for a wide range of applications.
7 Acknowledgement
This work is generously supported by NSF IIS-1949634, NSF SES-2128623, and has benefited from the Microsoft Accelerate Foundation Models Research (AFMR) grant program.
@@ A.1 Model input and output demonstration
Figure 6: Demo of interactive chat interface with the 3D-GRAND model.
A.2 Training Data
There are two flavors of models that we fine-tuned: one grounded object reference model, and one grounded QA model. The grounded object reference model was trained using the grounded object reference data on 3D-FRONT train split, which consist of 234,791 3D-text pairs, each of which are densely grounded. This model was used to generate the ScanRefer results presented in Table 4, 5, and Figure 4 The grounded QA model was trained using a subset of 200k grounded QA: object existence data from the 3D-FRONT train split. The reason that we select a subset of 200k QAs is simply because the entire grounded QA dataset is too large and we do not have enough resource to train on all data. However, as shown in Table 2 and Figure 4, we find even such a subset is very effective in reducing hallucinations in 3D-LLMs.
We provide official data splits of train, val and test (90%, 5%, 5%) in our dataset release. The val and test proportion might seem small, but given our dataset’s million-scale, they should be sufficiently large for any development and evaluation purposes.
A.3 Training Details
The two flavors of model mentioned above are LoRA-finetuned [28] based off Llama-2. We use DeepSpeed ZeRO-2 [55] and FlashAttention [16] to save GPU memory and speed up training. The model is trained in BF16 precision on 12 NVIDIA A40 GPUs with a combined batch size of 96 and a learning rate of 2e-4. We use the AdamW [45] optimizer with a weight decay of 0.01 and a cosine learning rate scheduler. We train the mode for 10k steps, which takes approximately 48 hours.
Appendix B Additional 3D-GRAND Data Collection
@@ D.1.2 Crowd-sourcing Validity
Before presenting any tasks to the workers, we present them with a set of instruction tasks that show an example annotation, the correct response (as determined by us), and the reason why that response is correct. They are paired with an incorrect example and an explanation of why it is incorrect in order to ensure unbiased annotations. Examples of qualifying instructions are shown in 11. These instructions are intentionally brief, as we found through trial-and-error that longer, paragraph-based instructions were largely ignored by annotators.